OpenStudy (anonymous):

let f(x)=sqr of x, g(x)=1/x-4 and h(x)= sqr of 5-x determine when the following functions are continuous. a) (fog)(x) b) (gof)(x)

OpenStudy (jdoe0001):

\(\bf f(x) = \sqrt{x}\qquad g(x) = \cfrac{1}{x-4}\qquad h(x)= \sqrt{5-x}\\\quad \\ (fog)(x) \implies f(\quad g(x)\quad )\implies \sqrt{\left(\cfrac{1}{x-4}\right)} \implies \cfrac{\sqrt{1}}{\sqrt{x-4}} \\\quad \\ (gof)(x) \implies g(\quad f(x)\quad )\implies \cfrac{1}{\left(\sqrt{x}\right)-4}\)